Consider the following situation after a government intervention. In this market, the quantity demanded is zero when the price is above 200 and quantity supplied is zero when the price is zero. Before the government intervention, the market equilibrium quantity and price were 100 and 10. After the government imposed a tax of $2 per unit, the price to consumers increased to $11 and only 90 units were

a. Calculate the consumer surplus after the government intervention.

b. Calculate the producer surplus after the government intervention.

c. Calculate the deadweight loss created by this policy
Answer #1

a) After the government intervention, the consumer surplus falls to area colored in Yellow whose area is (1/2) * (200 - 11) * 90 = 8505

b) After the government intervention, the producer surplus falls to area colored in Blue whose area is (1/2) * (9 - 0) * 90 = 405

c) The deadweight loss is the area colored in Orange + Green whose area is (1/2) * (11 - 9) * (100 - 90) = 10